# Permutation twisted cohomology, remixed

### v2

@@ 3.1. Removing contractible summands
Set $K_{i}:=kX_{l}$, by construction $d_{i}|_{K_{i}}$ is an isomorphism and maps canonical permutation basis to canonical permutation basis. We next modify the canonical permutation basis of $C_{n}$ to obtain a direct summand $D_{i}$ for which $d_{i}(D_{i})\subseteq D_{i-1}$. Set $C_{i}^{\prime}:=kX_{1}\oplus\cdots\oplus\widehat{kX}_{l}\oplus\cdots\oplus kX_{m}$, so we have $C_{i}=C_{i}^{\prime}\oplus K_{i}$. For each canonical permutation basis element $x$ of $C_{i}$ (i.e. some $x\in X_{j}$ for $j\neq l$) replace $x$ with $x^{\prime}:=x-((d_{i}|_{kX_{l}})^{-1}\circ p_{h}^{\prime}\circ d_{i})(x)$, where $p_{h}^{\prime}$ denotes projection onto $kY_{h}^{\prime}$ and $(d_{i}|_{kX_{l}})^{-1}$ denotes the inverse of the isomorphic projection of $kX_{l}$ onto $kY_{h}^{\prime}$. A straightforward computation shows $d_{i}(x^{\prime})\in D_{i-1}$. By construction, $((d_{i}|_{kX_{l}})^{-1}\circ p_{h}^{\prime}\circ d_{i})(x)\in K_{i}$, and the collection $X^{\prime}$ of the $x^{\prime}$ forms a $G$-set, with $X^{\prime}\cong X_{1}\sqcup\cdots\sqcup\widehat{X}_{l}\sqcup\cdots\sqcup X_{m}$. Set $D_{i}:=kX^{\prime}$, then we have a direct sum decomposition $C_{i}=D_{i}\oplus K_{i}$. Finally, for all $j\neq i,i-1$, set $D_{j}=C_{j}$. It follows that we have a direct sum decomposition $C=K\oplus D$. Finally, observe that for each $x^{\prime}$, under the image of $d_{i}$, the canonical permutation basis representation of $x^{\prime}$ in $D_{i-1}$ is identical to the canonical permutation basis representation of $x$ in $D_{i-1}$, therefore the stabilizers grow property holds for the differential $d_{i}$.
@@
Let $G$ be a finite $p$-group and let $C$ be an indecomposable effective endotrivial of $kG$-modules. Then $C$ arises from a chain complex $C^{?}$ of free $k\Gamma_{G}$-modules.
@@
Recall every effective endotrivial (after a possible shift) corresponds to a real representation $V$ of $\mathbb{R}G$-modules. The representation sphere $S^{V}$ is a $G$-CW-complex, therefore produces a (reduced) chain complex $C^{?}$ for which $C^{1}$ is an endotrivial with the same h-marks as $C$. By Corollary 3.11, we can remove all contractible summands from $C^{?}$ until $C^{1}$ is indecomposable, and since two indecomposable endotrivials with the same h-marks are isomorphic, the result follows. ∎
Moreover, we can completely identify the homomorphisms $k\to C$ for an effective endotrivial $C$.
@@
The existence of such a homomorphism $f$ is equivalent to the inclusion $T\subseteq\ker(d_{i}),$ and this inclusion holds since if not, Proposition 3.10 implies the existence of a contractible summand of $C^{1}$, hence a contractible summand of $C^{?}$, which cannot occur. The last statement is straightforward. ∎
@@ 4. An open cover of the Balmer spectrum via endotrivials
For the rest of the paper, we assume $G$ is a finite $p$-group.
4.1. $p$-local quasi-isomorphisms
@@
Set $C:=C_{V}$. First, $V$ is equivalently a faithful $\mathbb{R}[G/N]$-module, so it suffices to assume $N=1$ after replacing $G$ with $G/N$, and it suffices to show $C_{n}$ is indecomposable and projective. Since $C$ is effective, it follows by an inductive argument that for all subgroups $H\leq G$, $\Psi^{H}(C)$ is isomorphic to an indecomposable complex in $\mathscr{K}({G}/\!\!/{H})$ whose highest nonzero homological degree is $h_{C}(H)$ (see [Lin18, Proposition 5.8.11]). In particular, $n=h_{C}(1)$ is the maximal nonzero degree of $C$, since $C$ is indecomposable.
@@
Here, we have $C_{j}=M_{1}\oplus M_{2}$, $C_{j-1}=N_{1}\oplus N_{2}$, where $M_{1}$ and $N_{1}$ satisfy $\Psi^{H}(M_{1})=\Psi^{H}(C_{j})$ and $\Psi^{H}(N_{1})=\Psi^{H}(C)_{j-1}$ and $M_{2}$ and $N_{2}$ satisfy $\Psi^{H}(M_{2})=0$ and $\Psi^{H}(N_{2})=0$. With this setup, we have $\Psi^{H}(d_{1}^{1})=\Psi^{H}(d_{j})$. By applying [Mil24, Lemma 5.7] and its dual statement, $d_{1}^{1}$ is an isomorphism if and only if $\Psi^{K}(d_{1}^{1})$ is an isomorphism for all $K\geq H$. Since $\Psi^{H}(d_{1}^{1})=\Psi^{H}(d_{j})$ is an isomorphism, it follows by an inductive argument up the poset of subgroups $K$ of $G$ containing $H$ that $\Psi^{H}(d_{1}^{1})$ is an isomorphism. Now by a standard homological algebra argument (c.f. [BM23, Lemma 9.2]), $M_{1}\xrightarrow{d_{1}^{1}}N_{1}$ splits off as a contractible direct summand of $C$, a contradiction since $C$ was assumed to be indecomposable. Therefore, $j=h_{C}(H)$. We conclude the highest homological degree $i$ for which the permutation module $k[G/H]$ can occur as a direct summand of $C_{i}$ is $i=h_{C}(H).$ In particular, since $V$ is faithful, only projective modules can occur in $C_{n}$.
It remains to show $C_{n}$ is indecomposable. Suppose for contradiction $C_{n}\cong P_{1}\oplus P_{2}$ for projective $kG$-modules $P_{1},P_{2}$. Since $\dim_{k}H_{n}(C)=1$, either $d_{n}|_{P_{1}}$ or $d_{n}|_{P_{2}}$ is injective, hence split injective, and it follows that a contractible chain complex containing $P_{1}$ or $P_{2}$ splits off from $C$, contradicting indecomposability of $C$. Thus, $C_{n}$ is indecomposable projective, as desired. ∎
In the situation described in Proposition 4.1, $\ker(d_{n})$ is the unique submodule of $k[G/N]$ of $k$-dimension 1. We are now ready to generalize the $a_{N}$ and $b_{N}$ homomorphisms constructed in [BG25, Definition 12.3] for every effective endotrivial $C$ and ($p$-)subgroup $H\leq G$.
@@
unique up to scaling, such that $\Psi^{H}(\iota^{H}_{C})$ is an isomorphism $k[h_{C}(H)]\cong\Psi^{H}(C)$ in $\operatorname{D}_{b}(k[{G}/\!\!/{H}])$. We have $\iota^{H}_{C}=\iota^{{}^{g}H}_{C}$ for all $g\in G$. Moreover, the image of $\iota^{H}_{C}$ in $C_{h_{C}(H)}$ is contained in an indecomposable direct summand isomorphic to $k[G/K]$, for some subgroup $K\geq H$.
@@
Since $C$ is effective, we may assume (after possibly shifting $C$) by Corollary 3.12 that there exists an indecomposable chain complex of free $k\Gamma_{G}$-modules $C^{?}$ such that $C\cong C^{1}$. Denote the differentials of $C^{?}$ by $d_{i}^{?}$. By Proposition 4.1, the chain complex of permutation k[G//H]k[{G}/\!\!/{H}]-modules $C^{H}$, which by Proposition 3.5 is isomorphic in $\mathscr{K}({G}/\!\!/{H})$ to $\Psi^{H}(C)$, contains up to homotopy a unique indecomposable k[G//H]k[{G}/\!\!/{H}]-module $M$ in top homological degree $h_{C}(H)$, and $\dim_{k}\ker(d_{h_{C}(H)}^{H})=1$. It follows by the stabilizers grow condition that in the canonical permutation basis of $C^{?}$, $M$ corresponds to a unique direct summand in homological degree $h_{C}(H)$.
Therefore, we are in the situation of Proposition 3.7 after restricting $d_{i}^{?}$ to $M$. Choosing a nonzero $m\in\ker(d_{h_{C}(H)}^{H})$ (this choice is unique up to scaling), Proposition 3.7 asserts the existence of a $m^{\prime}\in\ker(d_{h_{C}(H)}^{1})$ generating a (unique) $k$-dimension one submodule of $C^{1}_{h_{C}(H)}$. We have a unique (up to scaling) nonzero homomorphism $\iota^{H}_{C}$ of chain complexes as follows.
@@
We have that $\Psi^{H}(\iota^{H}_{C})$ is by construction a quasi-isomorphism. The statement $\iota^{H}_{C}=\iota^{{}^{g}H}_{C}$ for all $g\in G$ follows since $c_{g}\circ\Psi^{H}=\Psi^{H^{\prime}}$ for any $g\in G$ satisfying $^{g}H=H^{\prime}$. The final statement follows from the final statement of Proposition 3.7. ∎
@@
One has to take care in how the maps $\iota^{H}_{C}$ are constructed; such maps that become quasi-isomorphisms locally are non-unique (not even up to choice of identification as in [BG25, Remark 12.5]). For instance, let $G=D_{16}$, $C$ be the endotrivial
$C:=kD_{16}\xrightarrow{d_{2}}k[D_{16}/H_{1}]\oplus k[D_{16}/H_{2}]\xrightarrow{d_{1}}k,$
where $H_{1}$ and $H_{2}$ are non-conjugate non-central subgroups of order 2, $d_{2}$ is induced by the assignment $g\mapsto(gH_{1},gH_{2})$, and $d_{1}$ is induced by augmentation homomorphisms. In this case, we could choose $\iota^{H_{1}}_{C}$ and $\iota^{H_{2}}_{C}$ to be the inclusion
@@
In this case, both $\hat{\Psi}^{H_{1}}(\iota)$ and $\hat{\Psi}^{H_{2}}(\iota)$ are isomorphisms, a desired property. However, from the construction in Proposition 3.7, it follows that $\iota^{H_{1}}_{C}$ and $\iota^{H_{2}}_{C}$ satisfy $\Psi^{H_{1}}(\iota^{H_{2}}_{C})=0$ and vice versa, since neither $H_{1}\geq_{G}H_{2}$ or vice versa.
Additionally, given two subgroups $K,H\leq G$, $\iota^{H}_{C}$ and $\iota^{K}_{C}$ may coincide. For instance, any shift of the tensor unit has $\iota^{H}_{C}=\iota^{G}_{C}$ for all subgroups $H\leq G$.
@@
Let $H,K$ be subgroups of $G$ and let $C$ be an effective endotrivial of $kG$-modules. Then $\iota^{H}_{C}=\iota^{K}_{C}$ if and only if there exists a subgroup $B$ of $G$ such that $h_{C}(B)=h_{C}(H)=h_{C}(K)$ and $H,K\leq_{G}B$.
@@
The forward implication is straightforward by the construction in Proposition 4.2 - in particular, we may choose $B$ to be the vertex of the permutation $kG$-submodule containing the image of $\iota^{H}_{C}$. Conversely, assume $h_{C}(H)=h_{C}(K)$. Suppose there exists a subgroup $B$ satisfying $h_{C}(B)=h_{C}(H)=h_{C}(K)$ and $H,K\leq_{G}B$. From the construction in Proposition 4.2, it follows that the image of $\iota^{B}_{C}$ is the unique minimal submodule of a transitive permutation module isomorphic to $k[G/B^{\prime}]$ for some $B^{\prime}$ containing $B$. It follows that $\iota^{H}_{C}$ and $\iota^{K}_{C}$ also have image contained in this permutation module, and hence also have image the unique minimal submodule of $k[G/B^{\prime}]$, as desired. ∎
4.2. The open cover
Construction 4.5.
With the notation of [BG25, Definition 12.3], given any normal subgroup $N$ of $G$ of index $p$, the maps $a_{N}$ and $b_{N}$ are examples of such morphisms (however, $c_{N}$ is not!). For instance, the endotrivial

### v3

@@ 3.1. Removing contractible summands
Set $K_{i}:=kX_{l}$, by construction $d_{i}|_{K_{i}}$ is an isomorphism and maps canonical permutation basis to canonical permutation basis. We next modify the canonical permutation basis of $C_{n}$ to obtain a direct summand $D_{i}$ for which $d_{i}(D_{i})\subseteq D_{i-1}$. Set $C_{i}^{\prime}:=kX_{1}\oplus\cdots\oplus\widehat{kX}_{l}\oplus\cdots\oplus kX_{m}$, so we have $C_{i}=C_{i}^{\prime}\oplus K_{i}$. For each canonical permutation basis element $x$ of $C_{i}$ (i.e., some $x\in X_{j}$ for $j\neq l$) replace $x$ with $x^{\prime}:=x-((d_{i}|_{kX_{l}})^{-1}\circ p_{h}^{\prime}\circ d_{i})(x)$, where $p_{h}^{\prime}$ denotes projection onto $kY_{h}^{\prime}$ and $(d_{i}|_{kX_{l}})^{-1}$ denotes the inverse of the isomorphic projection of $kX_{l}$ onto $kY_{h}^{\prime}$. A straightforward computation shows $d_{i}(x^{\prime})\in D_{i-1}$. By construction, $((d_{i}|_{kX_{l}})^{-1}\circ p_{h}^{\prime}\circ d_{i})(x)\in K_{i}$, and the collection $X^{\prime}$ of the $x^{\prime}$ forms a $G$-set, with $X^{\prime}\cong X_{1}\sqcup\cdots\sqcup\widehat{X}_{l}\sqcup\cdots\sqcup X_{m}$. Set $D_{i}:=kX^{\prime}$, then we have a direct sum decomposition $C_{i}=D_{i}\oplus K_{i}$. Finally, for all $j\neq i,i-1$, set $D_{j}=C_{j}$. We have a direct sum decomposition $C=K\oplus D$. Finally, observe that for each $x^{\prime}$, under the image of $d_{i}$, the canonical permutation basis representation of $x^{\prime}$ in $D_{i-1}$ is identical to the canonical permutation basis representation of $x$ in $D_{i-1}$, therefore the stabilizers grow property holds for the differential $d_{i}$.
@@
Corollary 3.12.
Let $G$ be a finite group. The tt-functor
$(-)^{1}:\operatorname{K}_{b}(\operatorname{proj}(k\Gamma_{G}))\to\mathscr{K}(G)$
induced by evaluation at the trivial subgroup $1$ of $G$ is conservative, i.e., detects vanishing of objects.
Proof.
3.11 implies the functor $(-)^{1}:\operatorname{K}_{b}(\operatorname{proj}(k\Gamma_{G}))\to\operatorname{K}_{b}(\operatorname{perm}(kG))$ is conservative, and the canonical inclusion $\operatorname{K}_{b}(\operatorname{perm}(kG))\to\operatorname{K}_{b}(\operatorname{perm}(kG))^{\natural}\cong\mathscr{K}(G)$ is conservative as well. ∎
@@
Theorem 3.13.
Let $G$ be a finite $p$-group and let $C$ be an indecomposable effective endotrivial of $kG$-modules. Then $C$ arises from a chain complex $C^{?}$ of free $k\Gamma_{G}$-modules. In particular, any effective endotrivial arises from a chain complex of free $k\Gamma_{G}$-modules.
Proof.
Recall every effective endotrivial (after a possible shift) corresponds to a real representation $V$ of $\mathbb{R}G$-modules. The representation sphere $S^{V}$ is a $G$-CW-complex, therefore produces a (reduced) chain complex $C^{?}$ for which $C^{1}$ is an endotrivial with the same h-marks as $C$. By 3.11, we can remove all contractible summands from $C^{?}$ until $C^{1}$ is indecomposable, and since two indecomposable endotrivials with the same h-marks are isomorphic, the result follows. The last statement follows easily by considering a direct sum decomposition of an effective endotrivial into its indecomposable endotrivial summand and contractible summand. ∎
We also prove a tightening of 2.8 for effective endotrivials. The examples in Example 2.7 demonstrate this, but it is not hard to find counterexamples of the proposition for $C$ non-effective.
@@
Let $G$ be a $p$-group and $C\in\operatorname{Pic}(\mathscr{K}(G))$ be an indecomposable effective endotrivial complex. Then $C_{h_{C}(G)}=k$.
@@
Without loss of generality we assume $h_{C}(G)=0$. By 2.8 we have $C_{0}$ has a direct summand isomorphic to $k$ and for all $i\neq 0$, $k$ is not a direct summand of $C_{i}$. Furthermore, since $C$ is effective, by a standard $p$-permutation argument (see [LIN18, Proposition 5.8.11]), $C_{i}=0$ for negative $i$. If $h_{C}\equiv 0$ then $C=k[0]$ so there is nothing to show. Otherwise, there exists a subgroup $H$ such that $h_{C}(K)>0$, and in particular $h_{C}(1)>0$.
Now assume for contradiction that $C_{0}$ has another indecomposable summand $M\cong k[G/K]$ with $K<G$. Since $h_{C}(1)>0$, $p_{M}\circ d_{1}$ is surjective, and because $C$ arises from a complex of free $k\Gamma_{G}$-modules, the fiber of $p_{M}\circ d_{1}$ in $C_{1}$ consists of indecomposable modules $k[G/K^{\prime}]$ with $K^{\prime}\leq K$. Here $p_{M}$ denotes projection onto $p_{M}$.
If $h_{C}(K)=0$, then monotonicity implies $h_{\operatorname{Res}^{G}_{K}}(C)\equiv 0$ so $\Psi^{K}(C)\cong k$. Therefore, there must be a summand $M^{\prime}$ of $C_{1}$ also isomorphic to $k[G/K]$ in the fiber of $p_{M}\circ d_{1}$ to “cancel” $\Psi^{K}(M)$ via homotopy equivalence. But by [LIN18, Proposition 5.10.7] the composition $p_{M}\circ d_{1}\circ i_{M^{\prime}}$ is an isomorphism if and only if $\Psi^{K}(\iota_{M^{\prime}}\circ d_{1}\circ p_{M})$ is. Here $i_{M^{\prime}}$ denotes the inclusion from $M^{\prime}$ into $C_{1}$. Therefore a contractible summand splits off (c.f. [BM26, Lemma 9.2]), contradicting indecomposability of $C$.
Therefore, $h_{C}(K)>0$ must hold. So $\Psi^{K}(d_{1})$ is surjective, and there must also be a summand $M^{\prime}$ of $C_{1}$ in the fiber of $p_{M}\circ d_{1}$ such that $\Psi^{K}(p_{M}\circ d_{1}\circ i_{M^{\prime}})$ is surjective. Now a similar argument as before yields a contractible summand of $C$, a contradiction. We’ve exhausted all cases, thus $C_{0}=k$.
This final corollary will be critical in the sequel: it shows that nonzero homomorphisms $k\to C[s]$ for $C$ indecomposable and arising from a complex of free $k\Gamma_{G}$ modules, can be drawn without concern. If $C$ is endotrivial, [GAL25] refers to such morphisms as global sections.
@@
The existence of such a homomorphism $f$ is equivalent to the inclusion $T\subseteq\ker(d_{i}),$ and this inclusion holds since if not, 3.10 implies the existence of a contractible summand of $C^{1}$, hence a contractible summand of $C^{?}$, which cannot occur. The last statement is straightforward. ∎
@@ 4. An open cover of the Balmer spectrum via endotrivials
Notation: For the rest of the paper, we assume $G$ is a finite $p$-group.
4.1. Constructing forerunners
We construct our “$p$-local quasi-isomorphisms,” the forerunners, for effective endotrivials. This first proposition establishes the uniqueness of their targets.
@@ 4.1. Constructing forerunners
Set $C:=C_{V}$. First, $V$ is equivalently a faithful $\mathbb{R}[G/N]$-module, so it suffices to assume $N=1$ after replacing $G$ with $G/N$, and it suffices to show $C_{n}$ is indecomposable and projective. Since $C$ is effective, it follows by an inductive argument (using [LIN18, Proposition 5.8.11]) that for all subgroups $H\leq G$, $\Psi^{H}(C)$ is isomorphic to an indecomposable complex in $\mathscr{K}({G}/\!\!/{H})$ whose highest nonzero homological degree is $h_{C}(H)$ . In particular, $n=h_{C}(1)$ is the maximal nonzero degree of $C$, since $C$ is indecomposable.
@@
Here, we have $C_{j}=M_{1}\oplus M_{2}$, $C_{j-1}=N_{1}\oplus N_{2}$, where $M_{1}$ and $N_{1}$ satisfy $\Psi^{H}(M_{1})=\Psi^{H}(C_{j})$ and $\Psi^{H}(N_{1})=\Psi^{H}(C)_{j-1}$ and $M_{2}$ and $N_{2}$ satisfy $\Psi^{H}(M_{2})=0$ and $\Psi^{H}(N_{2})=0$. With this setup, we have $\Psi^{H}(d_{1}^{1})=\Psi^{H}(d_{j})$. By applying [MIL24, Lemma 5.7] and its dual statement, $d_{1}^{1}$ is an isomorphism if and only if $\Psi^{K}(d_{1}^{1})$ is an isomorphism for all $K\geq H$. Since $\Psi^{H}(d_{1}^{1})=\Psi^{H}(d_{j})$ is an isomorphism, it follows by an inductive argument up the poset of subgroups $K$ of $G$ containing $H$ that $\Psi^{H}(d_{1}^{1})$ is an isomorphism. Now by a standard homological algebra argument (c.f. [BM26, Lemma 9.2]), $M_{1}\xrightarrow{d_{1}^{1}}N_{1}$ splits off as a contractible direct summand of $C$, a contradiction since $C$ was assumed to be indecomposable. Therefore, $j=h_{C}(H)$. We conclude the highest homological degree $i$ for which the permutation module $k[G/H]$ can occur as a direct summand of $C_{i}$ is $i=h_{C}(H).$ In particular, since $V$ is faithful, only projective modules can occur in $C_{n}$.
It remains to show $C_{n}$ is indecomposable. Suppose for contradiction $C_{n}\cong P_{1}\oplus P_{2}$ for projective $kG$-modules $P_{1},P_{2}$. Since $\dim_{k}H_{n}(C)=1$, either $d_{n}|_{P_{1}}$ or $d_{n}|_{P_{2}}$ is injective, hence split injective, so a contractible chain complex containing $P_{1}$ or $P_{2}$ splits off from $C$, contradicting indecomposability of $C$. Thus, $C_{n}$ is indecomposable projective, as desired. ∎
In the situation described in 4.1, $\ker(d_{n})$ is the unique submodule of $k[G/N]$ of $k$-dimension 1.
Remark 4.2.
We recall the morphisms $a_{N},b_{N},c_{N}$ constructed by Balmer–Gallauer, [BG25, Definition 12.3]. Recall that given $N\triangleleft G$ a subgroup of $G$ of index $p$, one has the endotrivial of length $2^{\prime}$ inflated from $C_{p}$
$u_{N}=k[G/N]\to k[G/N]\to k,$
inflated from a truncated periodic resolution of the trivial module $k$. The above example is in the setting of $p>2$. Then we have $2^{\prime}+1$ morphisms in $\mathscr{K}(G)$, $a_{N}\colon k\to u_{N}$, $b_{N}\colon k\to u_{N}[-2^{\prime}]$, and if $p>2$, $c_{N}\colon k\to u_{N}[-1]$. The morphism $a_{N}$ in degree 0 is the identity map $k\xrightarrow{\operatorname{id}}k$. The morphisms $b_{N}$ and $c_{N}$ (if it exists) are induced by the coaugmentation homomorphism $k\to k[G/N]$.
For all $H\leq G$, exactly one of $\Psi^{H}(a_{N})$ or $\Psi^{H}(b_{N})$ is a quasi-isomorphism, c.f. [BG25, Proposition 12.9]. On the other hand, $c_{N}$ is $\otimes$-nilpotent.
We now generalize these maps for all endotrivials, as announced by the introduction. The theorem statement does not capture all the properties of the maps, their construction in the proof will be relevant.
@@
such that $\Psi^{H}(\iota^{H}_{C})$ is an isomorphism $k[h_{C}(H)]\cong\Psi^{H}(C)$ in $\operatorname{D}_{b}(k[{G}/\!\!/{H}])$. We have $\iota^{H}_{C}=\iota^{{}^{g}H}_{C}$ for all $g\in G$. Moreover, the image of $\iota^{H}_{C}$ in $C_{h_{C}(H)}$ is contained in an indecomposable direct summand isomorphic to $k[G/K]$, for some subgroup $K\geq H$.
@@
Since $C$ is effective, we may assume (after possibly shifting $C$) by Theorem 3.13 that there exists an indecomposable chain complex of free $k\Gamma_{G}$-modules $C^{?}$ such that $C\cong C^{1}$. Denote the differentials of $C^{?}$ by $d_{i}^{?}$. By 4.1, the chain complex of permutation k[G//H]k[{G}/\!\!/{H}]-modules $C^{H}$, which by 3.5 is isomorphic in $\mathscr{K}({G}/\!\!/{H})$ to $\Psi^{H}(C)$, contains up to homotopy (after removing contractible summands) a unique indecomposable k[G//H]k[{G}/\!\!/{H}]-module $M$ in top homological degree $h_{C}(H)$, and $\dim_{k}\ker(d_{h_{C}(H)}^{H})=1$. By the stabilizers grow condition, in the canonical permutation basis of $C^{?}$, $M$ corresponds to a unique direct summand in homological degree $h_{C}(H)$.
Therefore, we are in the situation of 3.7 after restricting $d_{i}^{?}$ to $M$. Choosing a nonzero $m\in\ker(d_{h_{C}(H)}^{H})$, 3.7 asserts the existence of a $m^{\prime}\in\ker(d_{h_{C}(H)}^{1})$ generating a uniquely determined $k$-dimension one submodule of $C^{1}_{h_{C}(H)}$. We have a unique (up to scaling) nonzero homomorphism $\iota^{H}_{C}$ of chain complexes as follows:
@@
We have that $\Psi^{H}(\iota^{H}_{C})$ is by construction a quasi-isomorphism. The statement $\iota^{H}_{C}=\iota^{{}^{g}H}_{C}$ for all $g\in G$ follows since $c_{g}\circ\Psi^{H}=\Psi^{H^{\prime}}$ for any $g\in G$ satisfying $^{g}H=H^{\prime}$. The final statement follows from the final statement of 3.7. ∎
@@
We refer to the maps constructed in Theorem 4.3 as forerunners, and $\iota^{H}_{C}$ as the $H$-forerunner of $C$, as the maps herald the $p$-local isomorphisms $\hat{\Psi}^{H}(C)\cong k[h_{C}(H)]$. We have much to say about these maps.
  1. (a)
    One can also define these via the trace homomorphism
    $t^{G}_{H}:\operatorname{Hom}_{\mathscr{K}(H)}(\operatorname{Res}^{G}_{H}C,\operatorname{Res}^{G}_{H}D)\to\operatorname{Hom}_{\mathscr{K}(G)}(C,D)$
    defined by the assignment
    $f\mapsto\sum_{g\in[G/H]}g\cdot f(g^{-1}\cdot-).$
    See [BEN98a, Definition 3.6.2]. Setting $C=k$ and $D$ a shifted effective endotrivial, we are in the situation of Theorem 4.3, and have a morphism $f\colon k\to C^{H}[-h_{C}(G)]\hookrightarrow C^{1}[-h_{C}(H)]$. This morphism is $K$-stable for some maximum $K\geq H$, hence can be regarded as a morphism $f\in\operatorname{Hom}_{\mathscr{K}(K)}(k,\operatorname{Res}^{G}_{K}C[-h_{C}(H)])$. Setting $\iota^{H}_{C}=t^{G}_{K}(f)$ obtains our desired morphism.
  2. (b)
    The forerunner construction is quite ad-hoc, and appears to not coincide with any obvious topological construction, as the map from the tensor unit is a coaugmentation homomorphism. Moreover, their construction depends on a certain choice of direct sum decomposition; although this will not matter for our purposes, it still is dissatisfying, and in general the uniqueness (even up to scalar) of the forerunners is not clear. Whether a more natural construction, or a topological interpretation of these maps exist, are pertinent questions; we expect answers in the affirmative for both. Such a construction would allow for a much better understanding of these morphisms.
    For instance, with our current construction it is unclear whether the following natural property holds: given two effective endotrivials $C_{1},C_{2}$ and subgroup $H\leq G$, do we have an identification, up to scalars and contractibles, between the images of $\iota^{H}_{C_{1}\otimes C_{2}}$ and $\iota^{H}_{C_{1}}\iota^{H}_{C_{2}}$?
  3. (c)
    By 4.1 and Corollary 3.14, for any effective endotrivial $C$, $\iota^{1}_{C}$ and $\iota^{G}_{C}$ are unique up to scaling, since if $C$ is indecomposable, $C_{h_{C}(1)}$ is indecomposable and $C_{h_{C}(G)}=k$. It follows that $\iota^{H}_{C_{1}\otimes C_{2}}=\iota^{H}_{C_{1}}\iota^{H}_{C_{2}}$ up to scaling for $H=1$ and $H=G$.
  4. (d)
    One has to take care in how the forerunners are constructed, as such maps that become quasi-isomorphisms locally are non-unique (not even up to choice of identification as in [BG25, Remark 12.5]). For instance, let $G=D_{2^{n}}$ for $n\geq 3$, and $C$ be the endotrivial of Example 2.7.
    $C:=kD_{16}\xrightarrow{}k[D_{16}/H_{1}]\oplus k[D_{16}/H_{2}]\xrightarrow{}k,$
    where $H_{1}$ and $H_{2}$ are non-conjugate non-central subgroups of order 2 and the differentials are induced by augmentation homomorphisms. In this case, we could choose $\iota^{H_{1}}_{C}$ and $\iota^{H_{2}}_{C}$ to be the inclusion
@@
    In this case, both $\hat{\Psi}^{H_{1}}(\iota)$ and $\hat{\Psi}^{H_{2}}(\iota)$ are isomorphisms, a desired property. However, from the construction in 3.7, one observes that $\iota^{H_{1}}_{C}$ and $\iota^{H_{2}}_{C}$ satisfy $\Psi^{H_{1}}(\iota^{H_{2}}_{C})=0$ and vice versa, since neither $H_{1}\geq_{G}H_{2}$ or vice versa. This property is established in full generality in Lemma 4.11, and are necessary for our open cover to stratify the closed points of $\operatorname{Spc}(\mathscr{K}(G))$.
  5. (e)
    Given two non-conjugate subgroups $K,H\leq G$, $\iota^{H}_{C}$ and $\iota^{K}_{C}$ may coincide. As an easy example, any shift of the tensor unit has $\iota^{H}_{C}=\iota^{G}_{C}$ for all subgroups $H\leq G$.
@@
Let $H,K$ be subgroups of $G$ and let $C$ be an effective endotrivial of $kG$-modules. Then we may construct $\iota^{H}_{C}=\iota^{K}_{C}$ if and only if there exists a subgroup $B\leq G$ such that $h_{C}(B)=h_{C}(H)=h_{C}(K)$ and $H,K\leq_{G}B$.
@@
The forward implication follows from the construction in Theorem 4.3 - in particular, one may choose $B$ to be the vertex of the permutation $kG$-submodule containing the image of $\iota^{H}_{C}$. Conversely, assume $h_{C}(H)=h_{C}(K)$. Suppose there exists a subgroup $B$ satisfying $h_{C}(B)=h_{C}(H)=h_{C}(K)$ and $H,K\leq_{G}B$. From the construction of Theorem 4.3, the image of $\iota^{B}_{C}$ is the unique minimal submodule of a transitive permutation module isomorphic to $k[G/B^{\prime}]$ for some $B^{\prime}$ containing $B$. Then $\iota^{H}_{C}$ and $\iota^{K}_{C}$ $p$-locally also have image contained in this permutation module, and hence also have image the unique minimal submodule of $k[G/B^{\prime}]$, as desired. ∎
4.2. An open cover of the spectrum
Using the forerunners, we construct the promised open cover of $\operatorname{Spc}(\mathscr{K}(G))$. Recall that every prime of $\operatorname{Spc}(\mathscr{K}(G))$ is of the form $\mathscr{P}(H,\mathfrak{p}):=\hat{\psi}^{H}(\mathfrak{p})$ for some subgroup $H\leq G$ and $\mathfrak{p}\in V_{{G}/\!\!/{H}}:=\operatorname{Spc}(\operatorname{D}_{b}(kG))$, and $\mathscr{P}(H,\mathfrak{p})=\mathscr{P}(H^{\prime},\mathfrak{p}^{\prime})$ if and only if $H^{\prime}={}^{g}H$ and $\mathfrak{p}^{\prime}={}^{g}\mathfrak{p}$ for some $g\in G$.
Construction 4.6.
With the notation of [BG25, Definition 12.3], given any normal subgroup $N\leq G$ of index $p$, the maps $a_{N}$ and $b_{N}$ are examples of such morphisms (however, $c_{N}$ is not!). The endotrivial